# Communication-Efficient Stochastic Distributed Learning

### v1

@@ B-A Key bounds
$\delta$ $\coloneqq 1-\frac{|\tilde{\lambda}_{\max}|\rho\tau\gamma}{2},q_{4}\coloneqq\frac{4N\tau^{2}\gamma^{2}c_{0}+c_{3}}{1-\delta}$
$c_{0}$ $\coloneqq\gamma^{2}\tau\beta_{0},c_{1}\coloneqq\gamma^{3}\tau^{2}\beta_{1},c_{2}\coloneqq{12L^{2}\gamma^{2}}\|\widehat{\mathbf{V}}^{-1}\|^{2}N,$
$c_{3}$ $\coloneqq 4\gamma^{2}(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})\|\widehat{\mathbf{V}}^{-1}\|^{2}\tau^{2}N+{12L^{2}\gamma^{2}}\|\widehat{\mathbf{V}}^{-1}\|^{2}\tau^{2}N,$
$q_{1}$ $\coloneqq c_{0}\kappa_{1}\gamma\tau+c_{1},q_{2}\coloneqq c_{2},q_{3}\coloneqq 16\tau^{3}\gamma^{2}Nc_{0}.$
Proof.
When $\gamma\leq\min\{\bar{\gamma}_{1},\bar{\gamma}_{2}\}$, it follows that
$\gamma\tau\leq\frac{2}{|\tilde{\lambda}_{\max}|\rho},\quad|\tilde{\lambda}_{\min}|\rho\tau\gamma<2,\quad\gamma^{2}\tau(L^{2}+\rho^{2}d_{u}^{2})<\frac{1/16}{\tau-1}.$ (36)
Using (28) and Assumption 4, we have
$\|\widehat{\mathbf{h}}_{k}\|^{2}\leq c_{0}\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+c_{1}\|\widehat{\mathbf{d}}_{k}\|^{2}+c_{2}\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}+c_{3}\sigma^{2};$
using this fact together with (29) and (17), and by $\gamma<1$, we can then derive that (35) holds. ∎
B-B Theorem 1
We start our proof by recalling that the following inequality holds for and $L$-smooth function $f$, $\forall y,z\in\mathbb{R}^{n}$ [36]:
@@ B-B Theorem 1
Based on (30), substituting $y=\bar{x}_{k+1}$ and $z=\bar{x}_{k}$ into (37), using Assumption 4 and (27), we get
@@
$\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\gamma\mathbb{E}[\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}(\sum_{t}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)+\epsilon_{k})\rangle]$
$+\frac{\gamma^{2}L}{2}\mathbb{E}[\|\frac{1}{N}(\sum_{t}\sum_{i}g_{i}(\phi_{i,k}^{t})+\epsilon_{k})\|^{2}]$
$\leq\mathbb{E}[F(\bar{x}_{k})]-\gamma[\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}(\sum_{t}\sum_{i}\nabla f_{i}(\phi_{i,k}^{t})+\epsilon_{k})\rangle]$
$+2\gamma^{2}\tau L\mathbb{E}[\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]+{2\gamma^{2}\tau^{2}L\sigma^{2}}$
$+\frac{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau\mathbb{E}\left[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\tau\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2}\right]}{N}.$
@@
$-\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}\sum_{t}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)+\frac{1}{N}\epsilon_{k}\rangle$
$\leq-\sum_{t}\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}\sum_{i}\nabla f_{i}(\phi_{i,k}^{t})\rangle+\frac{\tau}{4}\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}$
$+\frac{1}{N^{2}\tau}\epsilon_{k}^{2}$
@@
$+\frac{\tau}{4}\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}+\frac{1}{N^{2}\tau}\epsilon_{k}^{2}$
$\leq-\frac{\tau}{4}\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}-\frac{1}{2}\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}$
$+(\frac{L^{2}}{2N}+\frac{2\rho^{2}d_{u}^{2}}{N})\|\widehat{\mathbf{\Phi}}_{k}\|^{2}+\frac{2\rho^{2}d_{u}^{2}\tau}{N}\|\mathbf{X}_{k}-\bar{\mathbf{X}}_{k}\|^{2}$
Now, combining the two equations above, and using (16), yields
$\mathbb{E}[F\left(\bar{x}_{k+1}\right)]\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\frac{\gamma\tau}{4}\mathbb{E}[\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}]$
$-\frac{\gamma}{2}(1-4\gamma\tau L)\sum_{t}\mathbb{E}[\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\left(\frac{\gamma L^{2}}{2N}+\frac{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2}}{N}\right)\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]$
$+\frac{2\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}\frac{18\gamma\rho^{2}d_{u}^{2}\tau+18\gamma^{2}\rho^{2}d_{u}^{2}\tau^{2}L}{N}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+{2\gamma^{2}\tau^{2}L\sigma^{2}}.$
@@
$-\frac{\gamma\tau}{4}\left(1-64\gamma\tau^{2}(\frac{\gamma L^{2}}{2}+2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2})\right)\cdot$
$\cdot\mathbb{E}[\|\nabla F\left(\bar{x}_{k}\right)\|^{2}]$
$-\frac{\gamma}{2}(1-4\gamma L\tau)\sum_{t}\mathbb{E}[\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\frac{1}{N}\left((\frac{72\gamma\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}\gamma^{2})\cdot\right.$
$\left.\cdot(\frac{\gamma L^{2}}{2}+{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2}})\right.$
$\left.+\frac{2\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}({18\gamma\rho^{2}d_{u}^{2}\tau+18\gamma^{2}\rho^{2}d_{u}^{2}\tau^{2}L})\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$
$+{2\gamma^{2}\tau^{2}L\sigma^{2}}+\left(\frac{\gamma L^{2}}{2}+{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2}}\right)4\tau^{2}\gamma^{2}\sigma^{2}.$
When $\gamma\leq\min\{\bar{\gamma}_{3},\bar{\gamma}_{4}\}$, then
$\gamma\tau^{2}(\frac{\gamma L^{2}}{2}+2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2})<\frac{1}{128},\quad\gamma L\tau<\frac{3}{16},$ (38)
@@
$\mathbb{E}[F\left(\bar{x}_{k+1}\right)]\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\frac{\gamma\tau}{8}\mathbb{E}[\|\nabla F\left(\bar{x}_{k}\right)\|^{2}]+$
@@
$+\frac{1}{N}\left((\frac{72\gamma\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}\gamma^{2})\cdot\right.$
$\left.\cdot(\frac{\gamma L^{2}}{2}+{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2}})\right.$
$\left.+\frac{2\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}({18\gamma\rho^{2}d_{u}^{2}\tau+18\gamma^{2}\rho^{2}d_{u}^{2}\tau^{2}L})\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$
$+{2\gamma^{2}\tau^{2}L\sigma^{2}}+\left(\frac{\gamma L^{2}}{2}+{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2}}\right)4\tau^{2}\gamma^{2}\sigma^{2}.$
Rearranging the above relation we get
@@
$+\frac{8}{N\gamma\tau}\left((\frac{72\gamma\tau^{2}}{|\tilde{\lambda}_{\max}|\rho}+(18+\frac{36\gamma\tau\rho^{2}d_{u}^{2}}{|\tilde{\lambda}_{\max}|\rho})16\tau^{3}\gamma^{2})\cdot\right.$
$\left.\cdot(\frac{\gamma L^{2}}{2}+{2\gamma^{2}L\rho^{2}d_{u}^{2}\tau+2\gamma\rho^{2}d_{u}^{2}})\right.$
$\left.+\frac{2\gamma\tau}{|\tilde{\lambda}_{\max}|\rho}({18\gamma\rho^{2}d_{u}^{2}\tau+18\gamma^{2}\rho^{2}d_{u}^{2}\tau^{2}L})\right)\cdot$
$\cdot\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+16\left({\gamma L^{2}}+{4\gamma^{2}L\rho^{2}d_{u}^{2}\tau+4\gamma\rho^{2}d_{u}^{2}}\right)\tau\gamma\sigma^{2}$
$+{16\gamma\tau L\sigma^{2}},$
@@
Summing over $k=0,1,\ldots,K-1$, using $-\tilde{F}\left(\bar{x}_{k}\right)\leq 0$ and $\gamma\leq 1$, it holds that
@@
where $c_{4}\coloneqq\gamma\kappa_{4}$, $c_{5}\coloneqq{16\gamma\tau L}+16\left({\gamma L^{2}}+{4\gamma^{2}L\rho^{2}d_{u}^{2}\tau+4\gamma\rho^{2}d_{u}^{2}}\right)\tau\gamma$.
@@
$\leq(\delta+\frac{q_{1}}{1-\delta})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+\frac{q_{2}}{1-\delta}\mathbb{E}[\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}]$ (40)
$+\frac{q_{3}}{1-\delta}\mathbb{E}[\left\|\nabla F(\bar{x}_{k})\right\|^{2}]+q_{4}\sigma^{2}\leq\bar{\delta}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+q_{4}\sigma^{2}+K\mathcal{D}_{k}$ (41)
@@
$\begin{split}R&\coloneqq\operatorname{\max}\{\frac{q_{3}}{1-\delta},\frac{q_{2}\tau^{2}}{1-\delta}\}.\end{split}$ (42)
Moreover, letting $\gamma\leq\min\{\bar{\gamma}_{1},\bar{\gamma}_{5}\}$, we have
$\bar{\delta}=\delta+\frac{q_{1}}{1-\delta}<1.$ (43)
@@
$\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]\leq\bar{\delta}_{k}\mathbb{E}[\|\widehat{\mathbf{d}}_{0}\|^{2}]+R\sum_{\ell=0}^{k-1}\bar{\delta}^{k-1-\ell}\mathcal{D}_{\ell}+\frac{q_{4}\sigma^{2}}{1-\bar{\delta}}$
@@
$\sum_{k=0}^{K-1}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]\leq\frac{\|\widehat{\mathbf{d}}_{0}\|^{2}}{1-\bar{\delta}}+\frac{R}{1-\bar{\delta}}\sum_{k=0}^{K-1}\mathcal{D}_{k}+\frac{q_{4}\sigma^{2}K}{1-\bar{\delta}}.$ (44)
@@
$\left(1-q_{5}\right)\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\frac{8\tilde{F}\left(\bar{x}^{0}\right)}{\gamma\tau}+q_{6}\|\hat{\mathbf{d}}_{0}\|^{2}+Kq_{7}\sigma^{2},$
@@
$q_{5}\coloneqq\frac{c_{4}R}{1-\bar{\delta}},\quad q_{6}\coloneqq\frac{c_{4}}{1-\bar{\delta}}\quad q_{7}\coloneqq\frac{c_{4}q_{4}}{1-\bar{\delta}}+c_{5}.$ (45)
When $\gamma\leq\frac{|\tilde{\lambda}_{\max}|^{2}\rho^{2}\tau^{2}}{8(\beta_{0}\kappa_{1}+\beta_{1})}$, we have $1-\bar{\delta}\geq\frac{|\tilde{\lambda}_{\max}|\rho\tau\gamma}{4},$ and $\frac{c_{4}}{1-\bar{\delta}}\leq\frac{4\kappa_{4}}{|\tilde{\lambda}_{\max}|\rho\tau}.$ Using this fact together with (42) and $\gamma\leq 1$, we have $q_{5}\leq\max\{\frac{128\kappa_{4}\tau^{2}N\beta_{0}\gamma}{|\tilde{\lambda}_{\max}|^{2}\rho^{2}\tau},\frac{24L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}N\gamma}{|\tilde{\lambda}_{\max}|^{2}\rho^{2}\tau^{2}}\}.$ Therefore, when $\gamma\leq\min\{1,\bar{\gamma}_{6},\bar{\gamma}_{7}\}$, we have
$q_{5}\leq\frac{1}{2},$ (46)
@@
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\frac{16\tilde{F}\left(\bar{x}^{0}\right)}{\gamma\tau K}+\frac{2q_{6}}{K}\|\hat{\mathbf{d}}_{0}\|^{2}+2q_{7}\sigma^{2}.$ (47)
By collecting all step-size conditions, if the step-size $\gamma<\bar{\gamma}_{\text{sgd}}\coloneqq\min_{i=1,2,\ldots,7}\bar{\gamma}_{i}$, then (36), (38), (43) and (46) hold, the states $\{\mathbf{X}_{k}\}$ generated by LT-ADMM converge to the neighborhood of the stationary point, concluding the proof.
Appendix C Convergence analysis for LT-ADMM-VR
C-A Key bounds
We start by deriving an upper bound for the variance of the gradient estimator $\mathbb{E}\left[\|g_{i}\left(\phi_{i,k}^{t}\right)-\nabla f_{i}\left({\phi}_{i,k}^{t}\right)\|^{2}\right]$. Define $t_{i}^{k}$ as the averaged consensus gap of the auxiliary variables of $\{r_{i,h,k}^{k}\}_{h=1}^{m_{i}}$ at node $i$:

### v2

@@ 8.1 Key bounds
$\delta$ $\coloneqq 1-\frac{\lambda_{l}\rho\tau\beta}{2},$
$\beta_{0}$ $\coloneqq\frac{72\beta\tau^{2}}{\lambda_{l}\rho}+144\tau^{3}\beta^{2}$
$c_{0}$ $\coloneqq\gamma^{2}(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})2\tau L^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2}\beta_{0},$
$c_{1}$ $\coloneqq\gamma^{4}\frac{6L^{2}}{\beta^{2}}N\|\widehat{\mathbf{V}}^{-1}\|^{2},$
$c_{2}$ $\coloneqq\gamma^{4}(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})L^{2}32\tau^{4}\|\widehat{\mathbf{V}}^{-1}\|^{2},$
$c_{3}$ $\coloneqq\gamma^{4}(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})8L^{2}N\tau^{3}\|\widehat{\mathbf{V}}^{-1}\|^{2}$
$+\gamma^{2}(1+2\rho^{2}\|\tilde{\mathbf{L}}\|^{2})2\tau^{2}N\|\widehat{\mathbf{V}}^{-1}\|^{2}+6L^{2}\frac{\gamma^{4}}{\beta^{2}}N\tau^{2}\|\widehat{\mathbf{V}}^{-1}\|^{2},$
Proof 8.7.
When $\beta<\frac{2}{\tau\lambda_{u}\rho}$ and $\gamma\leq\bar{\gamma}_{1}$, using (28), (29) and Assumption 4, we have
$\|\widehat{\mathbf{h}}_{k}\|^{2}$ $\leq c_{0}\|\widehat{\mathbf{d}}_{k}\|^{2}+c_{1}\|\sum_{t}\overline{\nabla F}(\boldsymbol{\Phi}^{t}_{k})\|^{2}$
$+c_{2}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}]+c_{3}\sigma^{2},$
together with (17) we can then derive that (37) holds.
8.2 Theorem 1
We start our proof by recalling that the following inequality holds for all $L$-smooth function $f$, $\forall y,z\in\mathbb{R}^{n}$ [38]:
@@ 8.2 Theorem 1
Based on (30), substituting $y=\bar{x}_{k+1}$ and $z=\bar{x}_{k}$ into (38), using Assumption 4, we get
@@
$\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\gamma\mathbb{E}[\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}\sum_{t}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\rangle]$
$+\frac{\gamma^{2}L}{2}\mathbb{E}[\|\frac{1}{N}\sum_{t}\sum_{i}g_{i}(\phi_{i,k}^{t})\|^{2}]$
$\leq\mathbb{E}[F(\bar{x}_{k})]-\gamma[\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}\sum_{t}\sum_{i}\nabla f_{i}(\phi_{i,k}^{t}))\rangle]$
$+\gamma^{2}\tau L\mathbb{E}[\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]+{\gamma^{2}\tau^{2}L\sigma^{2}}.$
@@
$-\langle\nabla F\left(\bar{x}_{k}\right),\frac{1}{N}\sum_{t}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\rangle$
@@
$\leq-\frac{\tau}{2}\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}-\frac{1}{2}\sum_{t}\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}+\frac{L^{2}}{2N}\|\widehat{\mathbf{\Phi}}_{k}\|^{2}.$
Now, combining the two equations above and using (16), yields
$\mathbb{E}[F\left(\bar{x}_{k+1}\right)]\leq\mathbb{E}[F\left(\bar{x}_{k}\right)]-\frac{\gamma\tau}{2}\mathbb{E}[\left\|\nabla F\left(\bar{x}_{k}\right)\right\|^{2}]$
$-\frac{\gamma}{2}(1-2\gamma\tau L)\sum_{t}\mathbb{E}[\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\frac{\gamma L^{2}}{2N}\mathbb{E}[\|\widehat{\mathbf{\Phi}}_{k}\|^{2}]+{\gamma^{2}\tau^{2}L\sigma^{2}}.$
@@
$-\frac{\gamma\tau}{2}\left(1-16L^{2}\tau^{2}\gamma^{2}\right)\mathbb{E}[\|\nabla F\left(\bar{x}_{k}\right)\|^{2}]$
$-\frac{\gamma}{2}(1-2\gamma L\tau)\sum_{t}\mathbb{E}[\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\frac{\gamma L^{2}}{2N}\left(\frac{72\beta\tau^{2}}{\lambda_{l}\rho}+144\tau^{3}\beta^{2}\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$
$+{\gamma^{2}\tau^{2}L\sigma^{2}}+2\tau^{2}\gamma^{3}\sigma^{2}L^{2}.$
When $\gamma\leq\min\{\bar{\gamma}_{2},\bar{\gamma}_{3}\}$, then
$16L^{2}\tau^{2}\gamma^{2}\leq\frac{3}{4},\quad 2\gamma L\tau\leq\frac{3}{4},$ (39)
@@
$+\frac{\gamma L^{2}}{2N}\left(\frac{72\beta\tau^{2}}{\lambda_{l}\rho}+144\tau^{3}\beta^{2}\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$
$+{\gamma^{2}\tau^{2}L\sigma^{2}}+2\tau^{2}\gamma^{3}\sigma^{2}L^{2}.$
Rearranging the above relation, we get
@@
$+\frac{8}{\gamma\tau}\frac{\gamma L^{2}}{2N}\left(\frac{72\beta\tau^{2}}{\lambda_{l}\rho}+144\tau^{3}\beta^{2}\right)\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]$
$+{8\gamma\tau L\sigma^{2}}+16\tau\gamma^{2}\sigma^{2}L^{2},$
@@
Summing over $k=0,1,\ldots,K-1$, using $\tilde{F}\left(\bar{x}_{k}\right)\leq 0$, it holds that
@@
where
$c_{4}$ $\coloneqq\frac{4L^{2}}{N}\left(\frac{72\beta\tau}{\lambda_{l}\rho}+144\tau^{2}\beta^{2}\right)$ (41)
[…]
@@
$\leq(\delta+\frac{c_{0}}{1-\delta})\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+\frac{c_{1}\tau}{1-\delta}\mathbb{E}[\|\frac{1}{N}\sum_{i}\nabla f_{i}\left(\phi_{i,k}^{t}\right)\|^{2}]$
$+\frac{c_{2}}{1-\delta}\mathbb{E}[\|\nabla F(\bar{x}_{k})\|^{2}]+\frac{c_{3}}{1-\delta}\sigma^{2}$
$\leq\bar{\delta}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]+\frac{c_{3}}{1-\delta}\sigma^{2}+R\mathcal{D}_{k}$ (42)
@@
$\begin{split}R&\coloneqq\operatorname{\max}\{\frac{c_{2}}{1-\delta},\frac{c_{1}\tau^{2}}{1-\delta}\}.\end{split}$ (43)
Moreover, letting $\gamma\leq\bar{\gamma}_{4}$ and $\frac{1}{\tau\lambda_{u}\rho}\leq\beta<\frac{2}{\tau\lambda_{u}\rho}$, we have
$\bar{\delta}=\delta+\frac{c_{0}}{1-\delta}<1-\frac{\lambda_{l}}{4\lambda_{u}}.$ (44)
@@
$\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]\leq\bar{\delta}^{k}\mathbb{E}[\|\widehat{\mathbf{d}}_{0}\|^{2}]+R\sum_{\ell=0}^{k-1}\bar{\delta}^{k-1-\ell}\mathcal{D}_{\ell}+\frac{c_{3}\sigma^{2}}{1-\bar{\delta}}$
@@
$\sum_{k=0}^{K-1}\mathbb{E}[\|\widehat{\mathbf{d}}_{k}\|^{2}]\leq\frac{\|\widehat{\mathbf{d}}_{0}\|^{2}}{1-\bar{\delta}}+\frac{R}{1-\bar{\delta}}\sum_{k=0}^{K-1}\mathcal{D}_{k}+\frac{c_{3}\sigma^{2}K}{1-\bar{\delta}}.$ (45)
@@
$\left(1-q_{0}\right)\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\frac{8\tilde{F}\left(\bar{x}^{0}\right)}{\gamma\tau}+q_{1}\|\hat{\mathbf{d}}_{0}\|^{2}+Kq_{2}\sigma^{2},$
@@
$q_{0}\coloneqq\frac{c_{4}R}{1-\bar{\delta}},\quad q_{1}\coloneqq\frac{c_{4}}{1-\bar{\delta}}\quad q_{2}\coloneqq\frac{c_{4}c_{3}}{1-\bar{\delta}}+c_{5}.$ (46)
Since $1-\bar{\delta}\geq\frac{\lambda_{l}}{4\lambda_{u}}$ and $1-{\delta}\geq\frac{\lambda_{l}}{2\lambda_{u}}$, when $\gamma\leq\min\{1,\bar{\gamma}_{5},\bar{\gamma}_{6}\}$, we have
$q_{0}\leq\frac{1}{2},$ (47)
@@
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{D}_{k}\leq\frac{16\tilde{F}\left(\bar{x}^{0}\right)}{\gamma\tau K}+\frac{2q_{1}}{K}\|\hat{\mathbf{d}}_{0}\|^{2}+2q_{2}\sigma^{2}.$ (48)
By collecting all step-size conditions, if the step-size $\gamma<\bar{\gamma}_{\text{sgd}}\coloneqq\min_{i=1,2,\ldots,6}\bar{\gamma}_{i}$, then (48) holds, the states $\{\mathbf{X}_{k}\}$ generated by LT-ADMM converge to the neighborhood of the stationary point, concluding the proof.
9 Convergence analysis for LT-ADMM-VR
9.1 Key bounds
We start by deriving an upper bound for the variance of the gradient estimator $\mathbb{E}[\|g_{i}(\phi_{i,k}^{t})-\nabla f_{i}({\phi}_{i,k}^{t})\|^{2}]$. Define $t_{i}^{k}$ as the averaged consensus gap of the auxiliary variables of $\{r_{i,h,k}^{k}\}_{h=1}^{m_{i}}$ at node $i$: